Feather Friends, Inc., distributes a high-quality wooden birdhouse that sells for $80 per unit. Variable expenses are $40.00 per unit, and fixed expenses total $180,000 per year. Its operating results for last year were as follows:

Sales $ 2,160,000
Variable expenses 1,080,000
Contribution margin 1,080,000
Fixed expenses 180,000
Net operating income $ 900,000

Required:

Answer each question independently based on the original data:

1. What is the product's CM ratio?

2. Use the CM ratio to determine the break-even point in dollar sales.

3. If this year's sales increase by $59,000 and fixed expenses do not change, how much will net operating income increase?

4-a. What is the degree of operating leverage based on last year's sales?

4-b. Assume the president expects this year's sales to increase by 13%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

5. The sales manager is convinced that a 13% reduction in the selling price, combined with a $68,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $900,000 net operating income as last year?
Answer #1

1. What is the product's CM ratio?

Contribution margin per unit= Sales price per unit-Variable expenses per unit

= $80-40= $40

CM ratio= Contribution margin*100/Sales

= $40*100/80= 50%

2. Use the CM ratio to determine the break-even point in dollar sales.

Break even point in dollar sales= Fixed cost/CM ratio

= $180000/50%= $360000

3. If this year's sales increase by $59,000 and fixed expenses do not change, how much will net operating income increase?

Net operating income increases by= Increase in sales*CM ratio

= $59000*50%= $29500

4-a. What is the degree of operating leverage based on last year's sales?

Degree of operating leverage= Contribution margin/Net operating income

= $1080000/900000= 1.2

4-b. Assume the president expects this year's sales to increase by 13%. Using the degree of operating leverage from last year, what percentage increase in net operating income will the company realize this year?

Degree of operating leverage= Percentage change in net operating income/Percentage change in sales

1.2= Percentage change in net operating income/13%

Percentage change in net operating income= 13%*1.2= 15.6%

5. The sales manager is convinced that a 13% reduction in the selling price, combined with a $68,000 increase in advertising, would increase this year's unit sales by 25%.

a. If the sales manager is right, what would be this year's net operating income if his ideas are implemented?

New sales price= $80*87%= $69.6

New fixed expenses= $180000+68000= $248000

Increase in sales units= ($2160000/80)*1.25= 33750 units

Sales (33750*$69.6) $2349000
Variable expenses (33750*$40) 1350000
Contribution margin 999000
Fixed expenses 248000
Net operating income $751000

b. If the sales manager's ideas are implemented, how much will net operating income increase or decrease over last year?

Net operating income increases (decreases)= $751000-900000= $(149000)

6. The president does not want to change the selling price. Instead, he wants to increase the sales commission by $1.80 per unit. He thinks that this move, combined with some increase in advertising, would increase this year's sales by 25%. How much could the president increase this year's advertising expense and still earn the same $900,000 net operating income as last year?

New sales units= ($2160000/80)*1.25=  33750 units

Profit= Contribution- Fixed cost

$900000= ($80-40-1.80)*33750- Fixed cost

$900000= 1289250- Fixed cost

Fixed cost= $1289250-900000= $389250

Increase in advertising= $389250-180000= $209250